OpenStudy (anonymous):

how do i find the prime factorization of 425

jimthompson5910 (jim_thompson5910):

Divide 425 by 5 to get 425/5 = 85 Divide 85 by 5 to get 85/5 = 17 So 425 = 5*5*17